# SEED-Story: Multimodal Long Story Generation with Large Language Model

### v1

@@ 4.2 Key Features
Unlike existing story generation datasets, such as Flintstones and Pororo, which offer images at a resolution of 128x128, our story stream dataset provides high-resolution images of 480x768.
@@
Figure 5: Story visualization comparison between SEED-Story and baseline models. SEED-Story generates images with higher quality and better consistency.
@@ 5 Experiment
Table 2: Quantitative evaluation for story visualization.
[…]
LDM 67.29 0.7585
StoryGen 73.74 0.7573
SEED-Story 67.01 0.7793
5.1 Story Visualization
Previous story generation approaches primarily utilize diffusion models, focusing on visualizing story images. These models take the previous image and text as input, and then generate only the next image based on the current text prompt. For a fair comparison, we adapt our model to a visualization-only format. For StoryGen [27], we also train it to produce images with previous images and texts. For LDM [43], we only give it text-image pairs. The visual results are shown in Figure 5. SEED-Story model shows better style and character consistency and higher quality compared to baselines. We also conduct a quantitative evaluation in Table 2 to demonstrate our effectiveness.
5.2 Multimodal Story Generation
For comparative analysis, there are relatively few established methods for generating multimodal stories. To establish a baseline for comparison, we fine-tune the recently developed MM-interleaved model on our dataset. We detail the comparative results in Figure 6. The FID is employed to assess the visual quality of the generated images. Additionally, we ask GPT-4V to compare and choose a preferred option between each of the generation results of MM-interleaved and SEED-Story across several dimensions: Style Consistency, which evaluates the stylistic uniformity across different images; Story Engagement, which measures the ability of narratives to captivate and maintain audience interest; and Image-Text Coherence, which assesses the alignment and relevance between images and their accompanying texts. The appendix presents our visual results and details about prompt engineering methods.
For qualitative presentation, we demonstrate our SEED-story effectively helps users control the story plot. Figure 7 shows that using the same starting image but different initial texts, the narrative branches into two distinct storylines.
We provide more cases in Figure 8 and Figure 9 to prove our multimodal long story generation capabilities. SEED-story can generate long sequences with engaging plots and vivid images.
Figure 7: Examples of multimodal story generation from SEED-Story. It shows two narrative branches generated from the same initial image. The top branch starts with text referencing “the man in the yellow hat,” leading to images that include the character. The bottom branch starts without mentioning the man, resulting in stories that diverge from the first by excluding him.
Figure 8: Multimodal long story generation results of SEED-Story.
Figure 9: Multimodal long story generation results of SEED-Story.
Figure 10: The visualization of generating long stories with different attention mechanisms. The models adopting other mechanisms collapse when producing long sequence, while our model employing multimodal attention sink consistently generates high-quality images.
@@ 5.2 Multimodal Story Generation
5.3 Long Story Generation
To verify our effectiveness in long story generation, we conduct an experiment visualizing a long story using the SEED-Story model, but with varying attention mechanisms. We chunk our data into stories of length of 10 considering the training efficiency. We set the window size as the same as the training length. Qualitative results depict in Figure 10 demonstrate that window attention quickly collapses when the inference length exceeds the training length. Both dense attention and attention sink approaches fare better, yet still fail to produce meaningful images as the inference sequence lengthens. In contrast, the multimodal attention sink consistently produces high-quality images. In terms of efficiency, the multimodal attention sink exhibits significant improvement over dense attention, with only a modest increase in time and memory costs compared to window attention and vanilla attention sink. These additional costs stem from retaining extra image tokens in the KV cache. Quantitative results presented in Table 3 substantiate the above conclusion.
@@ 6 Conclusion
Appendix A Implementation Details
A.1 Visual Tokenization and De-tokenization
@@ A.3 De-tokenizer Adaptation
Appendix B Attention Map Visualization
@@
In this section, we present additional visualizations of attention maps. These maps are derived from various model runs, including varying data lengths, attention heads, and layers. The visualizations consistently reveal a pattern of attention focused on ’0’ tokens, punctuation, tokens adjacent to Begin-of-Image (BoI), and tokens adjacent to End-of-Image (EoI).
@@
Figure 11: Visualization of attention maps from various model runs, showcasing attention patterns across different data lengths, attention heads, and model layers. Notably, the maps highlight consistent focus on ’0’ tokens, punctuation, tokens adjacent to Begin-of-Image (BoI), and tokens adjacent to End-of-Image (EoI).
Appendix C Data Usage and License
C.1 Curious George
Curious George is released on PBS KIDS [37, 38], a not-for-profit institution. It is a production of Imagine, WGBH and Universal. Curious George and related characters, created by Margret and H.A. Rey, are copyrighted and trademarked by Houghton Mifflin Harcourt and used under license. Licensed by Universal Studios Licensing LLC. Television Series: ©2024 Universal Studios. The terms of use of them are provided in https://www.pbs.org/about/about-pbs/terms-of-use/.
C.2 Rabbids Invasion
“Rabbids Invasion” is a French-American computer-animated TV series that breathes life into the zany antics of Ubisoft’s popular Rabbids video game characters. Created by Jean-Louis Momus and featuring the voice of Damien Laquet, the show is a dynamic blend of humor and adventure tailored for a family audience. Since its debut on August 3, 2013, on France 3, the series has enjoyed multiple seasons and a global reach. The Rabbids are mischievous rabbit-like creatures whose escapades lead them into all sorts of unpredictable and hilarious situations, making “Rabbids Invasion” a delight for both kids and adults alike. Thanks to their release, we derive some subsets from the cartoon series Rabbids Invasion [4, 3].
C.3 The Land Before Time
The Land Before Time, an iconic animated film series created by Judy Freudberg and Tony Geiss and distributed by Universal Pictures, debuted in 1988 with significant contributions from Don Bluth, George Lucas, and Steven Spielberg. This franchise, consisting of an initial film followed by 13 sequels, a TV series, video games, and extensive merchandising, explores the adventures of five young dinosaurs who learn key life lessons about friendship and teamwork through their prehistoric trials. Despite the absence of the original creators in the sequels, the series has continued to captivate audiences, emphasizing themes of community and perseverance across its extensive narrative arc. Thanks to their release, we derive some subsets from their websites [51, 50].
C.4 Appreciation
Appendix D Story Visualization Comparison
In this section, we present additional visualization comparison of our SEED-Story and other story visualization methods. SEED-Story shows better image consistency and higher qualit, as shown in Figure 12.
Figure 12: More story visualization comparison of SEED-Story and other story visualization methods.
Appendix E Multimodal Story Generation Results
In this section, we present our multimodal story generation results of our SEED-Story. It keeps produce story image and text with high quality. Please see Figure 13.
We also showcase the visualization result of our model on Rabbids Invasion and The Land Before Time. Please see Figure 14 and 15.
Figure 13: Multimodal story generation results of SEED-Story.
Figure 14: Story visualization result on Rabbids Invasion data.
@@ Appendix E Multimodal Story Generation Results
Figure 15: Story visualization result on The Land Before Time data.
Appendix F Details about GPT-4V Evaluation
F.1 Comparative Evaluation
@@
To evaluate the effectiveness of MM-interleaved and SEED-Story in multimodal story generation, we initiate an experiment where each model produces a story of five segments, based on a common starting image and text. The segment limit is set to five to accommodate the constraints of GPT-4V, which can handle a maximum of ten images per input session. In total, we generate 180 stories for assessment. For evaluation, we employ GPT-4 or GPT-4V to determine which model produces the better story in each case, based on the framework established in L-Eval [2]. We calculate the win rate for each model to determine its performance relative to its counterpart. The prompt we used is shown below.
@@ F.2 Score Evaluation
Appendix G Story Video
@@
Appendix H Broader Impacts

### v2

@@ 4.2 Key Features
Unlike existing story generation datasets which offer images at a resolution of 128x128, our story stream dataset provides high-resolution images of 480x768.
@@
Figure 7: Examples of multimodal story generation from SEED-Story with the same initial image but different texts as the beginning. The top branch starts with the text referencing “the man in the yellow hat”, which leads to the generation of images that include the character. The bottom branch starts without mentioning the man, resulting in a different storyline.
@@ 5 Experiment
5.1 Quantitative Results
For quantitative evaluation of multimodal story generation, since there are relatively few established methods for generating multimodal stories, we first fine-tune the recently developed MM-interleaved model on our training dataset for a fair comparison as the baseline model. The quantitative results are listed in Figure 6. The FID is employed to assess the visual quality of the generated images. Additionally, we ask GPT-4V (“gpt-4-turbo-2024-04-09”) to compare and choose a preferred option between each of the generation results of MM-interleaved and SEED-Story across several dimensions: (a) Style Consistency, which evaluates the stylistic uniformity across different images; (b) Story Engagement, which measures the ability of narratives to captivate and maintain audience interest; (c) Image-Text Coherence, which assesses the alignment and relevance between images and their accompanying texts. The details of evaluation are introduced in Section G of Appendix. We also conduct a user study, which compares SEED-Story and MM-Interleaved in Section A of Appendix. The quantitative evaluations demonstrate that SEED-Story outperforms the baseline model in terms of story engagement and image-text coherence, and achieves slightly higher preference in image style consistency. We also provide quantitative comparisons of story visualization in Section E of Appendix.
Figure 8: The visualization of generating long stories with different attention mechanisms. Without multimodal attention sink, MLLM cannot generate meaningful long image sequences. As highlighted in the red boxes, other methods produce meaningless images in the later frames.
@@ 5.1 Quantitative Results
5.2 Qualitative Results
For qualitative evaluations, we first demonstrate that our SEED-story can effectively generate stories with different plots and corresponding illustrations based on the different beginnings provided by the user, as shown in Fig. 7. We provide more qualitative results of multimodal story generation in Section F of Appendix. As shown in Figure F, Figure G and Figure H, SEED-story can generate long sequences with engaging plots and vivid images. We further provide qualitative comparisons of story visualization in Section E of Appendix.
5.3 Multimodal Attention Sink
To verify the effectiveness of multimodal attention sink in long story generation, we conduct an experiment visualizing a long story using the SEED-Story model, but with varying attention mechanisms. We chunk our data into stories of length of 10 considering the training efficiency. We set the window size as the same as the training length. Qualitative results presented in Figure 8 demonstrate that window attention quickly collapses when the inference length exceeds the training length. Both dense attention and attention sink approaches fare better, yet still fail to produce meaningful images as the inference sequence lengthens. In contrast, the multimodal attention sink consistently produces high-quality images. In terms of efficiency, the multimodal attention sink exhibits significant improvement over dense attention, with only a modest increase in time and memory costs compared to window attention and vanilla attention sink. These additional costs stem from retaining extra image tokens in the KV cache. Quantitative results presented in Table 2 substantiate the above conclusion.
@@ 6 Conclusion
Appendix A User study
Criteria MM-Interleaved SEED-Story Ties
Image Quality 21.3% 66.2% 12.5%
Image Style Consistency 7.5% 78.8% 13.7%
Image Diversity 21.3% 75.0% 3.7%
Story Engagement 37.5% 55.0% 7.5%
Image-Story Coherence 5.0% 86.3% 8.7%
Table 3: The results of user study between MM-Interleaved, SEED-Story, and Ties
Participants were asked to choose their preference based on image quality, image style consistency, image diversity, story engagement, and image-story coherence. The results were obtained from 125 samples, as shown in Table 3.
The results indicate that SEED-Story clearly outperforms the baseline in image generation ability and text-to-image coherence. Additionally, SEED-Story shows a slightly higher preference in text quality for story generation.
Appendix B Ablation study on de-tokenizer adaptation
We find that the generated images before the de-tokenizer adaptation stage exhibit semantic relevance with consistent backgrounds and characters, thanks to MLLM’s context preservation, as shown in Figure A. However, they suffer from texture distortion and inconsistency in style. After de-tokenizer adaptation, the images show improved consistency in style and character appearance. The calculated FID scores in table 4 confirm that de-tokenizer adaptation enhances image quality.
Figure A: Generated story images with and without the 3rd stage: de-tokenizer adaptation. The images generated before the third stage preserve semantic information, with mostly correct backgrounds and characters. However, they display low-quality textures and inconsistency in cartoon style. Our third stage effectively enhances these aspects.
Model FID
Before 3rd stage 153.93
After 3rd stage 99.79
Table 4: FID scores before and after the 3rd stage.
Appendix C Implementation Details
C.1 Visual Tokenization and De-tokenization
@@ C.3 De-tokenizer Adaptation
Appendix D Analysis of Multimodal Attention Sink
D.1 Attention Map Visualization
@@
In this section, we present additional visualizations of attention maps. These maps are derived from various model runs, including varying data lengths, attention heads, and layers. The visualizations consistently reveal a pattern of attention focused on “0” tokens, punctuation, tokens adjacent to Begin-of-Image (BoI), and tokens adjacent to End-of-Image (EoI).
@@
Figure B: Visualization of attention maps from various model runs, showcasing attention patterns across different data lengths, attention heads, and model layers. Notably, the maps highlight consistent focus on ’0’ tokens, punctuation, tokens adjacent to Begin-of-Image (BoI), and tokens adjacent to End-of-Image (EoI).
D.2 Statistics on Token’s Attention
To demonstrate that more attention is paid to image tokens adjacent to <BOI> and <EOI>, we analyzed over 5600 attention maps from various models, layers, and input sequences, to identify “key tokens” with the highest mean attention values. For each attention map, we computed the mean attention value across the query dimension for every key and selected the top 10 keys. We then aggregated these results to determine how often each token appeared as a key token. Table 5 shows the tokens with the highest occurrences, supporting our multimodal attention sink mechanism, with most queries focusing on four key token categories.
  • Starting tokens (BOS)
  • Punctuation (“,” “.” and “;” …)
  • Image tokens near BOI (BOI, IMG04)
  • Image tokens near EOI (from IMG57 to IMG63, added EOI)
BOS IMG57 EOI IMG04 , IMG60
Key Token Occurrence 4320 4320 4320 4140 4140 4120
IMG61 IMG59 IMG62 IMG63 BOI IMG56
Key Token Occurrence 3730 3132 1651 607 603 361
Table 5: Key Token Occurrence for Various Tokens
Appendix E Story Visualization Comparison
Table 6: Quantitative evaluation for story visualization.
[…]
LDM 67.29 0.7585
StoryGen 73.74 0.7573
SEED-Story 67.01 0.7793
Previous story generation approaches primarily utilize diffusion models, focusing on visualizing story images. These models take the previous image and text as input, and then generate only the next image based on the current text prompt. For a fair comparison, we adapt our model to a visualization-only format. For StoryGen Liu et al. (2023a), we also train it to produce images with previous images and texts. For LDM Rombach et al. (2022), we only give it text-image pairs. The visual results are shown in Figure C. SEED-Story model shows better style and character consistency and higher quality compared to baselines. We also showcase the visualization result of our model on Rabbids Invasion and The Land Before Time. Please see Figure D and E. We further conduct a quantitative evaluation in Table 6 to demonstrate our effectiveness.
Figure C: Story visualization comparison of SEED-Story and other story visualization methods.
Figure D: Story visualization result on Rabbids Invasion data.
@@ Appendix E Story Visualization Comparison
Figure E: Story visualization result on The Land Before Time data.
Appendix F Multimodal Story Generation Results
In this section, we present more multimodal story generation results of our SEED-Story. It keeps produce story image and text with high quality. Figure F, Figure G, and Figure H prove our multimodal long story generation capabilities. SEED-story can generate long sequences with engaging plots and vivid images.
Figure F: Multimodal long story generation results of SEED-Story.
Figure G: Multimodal long story generation results of SEED-Story.
Figure H: Multimodal story generation results of SEED-Story.
Appendix G Details about GPT-4V Evaluation
G.1 Comparative Evaluation
@@
To evaluate the effectiveness of MM-interleaved and SEED-Story in multimodal story generation, we initiate an experiment where each model produces a story of five segments, based on a common starting image and text. The segment limit is set to five to accommodate the constraints of GPT-4V, which can handle a maximum of ten images per input session. In total, we generate 180 stories for assessment. For evaluation, we employ GPT-4 or GPT-4V to determine which model produces the better story in each case, based on the framework established in L-Eval An et al. (2023). We calculate the win rate for each model to determine its performance relative to its counterpart. The prompt we used is shown below.
@@ G.2 Score Evaluation
Appendix H Story Video
@@
Appendix I Data Usage and License
I.1 Curious George
Curious George is released on PBS KIDS PBS Kids (2024a; b), a not-for-profit institution. It is a production of Imagine, WGBH and Universal. Curious George and related characters, created by Margret and H.A. Rey, are copyrighted and trademarked by Houghton Mifflin Harcourt and used under license. Licensed by Universal Studios Licensing LLC. Television Series: ©2024 Universal Studios. The terms of use of them are provided in https://www.pbs.org/about/about-pbs/terms-of-use/.
I.2 Rabbids Invasion
“Rabbids Invasion” is a French-American computer-animated TV series that breathes life into the zany antics of Ubisoft’s popular Rabbids video game characters. Created by Jean-Louis Momus and featuring the voice of Damien Laquet, the show is a dynamic blend of humor and adventure tailored for a family audience. Since its debut on August 3, 2013, on France 3, the series has enjoyed multiple seasons and a global reach. The Rabbids are mischievous rabbit-like creatures whose escapades lead them into all sorts of unpredictable and hilarious situations, making “Rabbids Invasion” a delight for both kids and adults alike. Thanks to their release, we derive some subsets from the cartoon series Rabbids Invasion Animaj (2024b; a).
I.3 The Land Before Time
The Land Before Time, an iconic animated film series created by Judy Freudberg and Tony Geiss and distributed by Universal Pictures, debuted in 1988 with significant contributions from Don Bluth, George Lucas, and Steven Spielberg. This franchise, consisting of an initial film followed by 13 sequels, a TV series, video games, and extensive merchandising, explores the adventures of five young dinosaurs who learn key life lessons about friendship and teamwork through their prehistoric trials. Despite the absence of the original creators in the sequels, the series has continued to captivate audiences, emphasizing themes of community and perseverance across its extensive narrative arc. Thanks to their release, we derive some subsets from their websites TheLandBeforeTime (2024a; b).
I.4 Appreciation
Appendix J Broader Impacts